Nitrogen in a wastewater treatment plant is in the form of ammonia and ammonium ion and has a total ammonia concentration of 7.1 x 10^-4 M at pH = 7. Total ammonia concentration = [NH3] + [NH4^+]. NH3 partial pressure = 5 x 10^-10 atm. KH at 25-degree Celcius is 57 mol/(L x atm). Ka at 25-degree Celcius = 10^-9.26.

A. Calculate NH3 and [NH4+] in M at the present condition where pH = 7 at 25 degrees Celcius.

B. If the wastewater pH is raised to 9, calculate equilibrium [NH3] and [NH4] in M.
Answer #1

14-7= 7. a) -4-74 to =4574 10 7= 4-74+ los (NH] = p3+ lo CNk] (NHa) +[Nts] =7.|X 7.18167 E 181-97 %3D 3-88x15€m W. 182-97